OpenStudy (anonymous):

find an equation of the line containing the given pair of points {1/6,-1/3} and {5/6,3)

OpenStudy (anonymous):

Still here? I can assist you with this in like 5 minutes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Well, while I'm waiting for the person to answer my other question, what would you do to find the slope of the two points?

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

Ok. The slope formula is (y1 - y2 )/(x1 -x2). So for now, tell me what our values for each one is. As in what is my y1, my y2, blah blah

OpenStudy (anonymous):

1y 1x

OpenStudy (anonymous):

the equation will look like y = mx + b where m and b will be defined. to find m : (y2-y1)/(x2-x1) to find b : value of y when x = 0 i can help further, but i rather you use your current knowledge to proceed from this :)

OpenStudy (anonymous):

Well actually, the x1 y1 means "the x value of my first point, and the y value of my first point"

OpenStudy (anonymous):

So try that one again. ^_^ And be sure to also tell me what our x2,y2 value is. Hint: remember that the x value is the first number in a point, and the y value is the second value.

OpenStudy (anonymous):

This is my first night doing these, I really have no experience

OpenStudy (anonymous):

Okie dokie. Your points were: (1/6,-1/3) and (5/6,3) Or, even more pretty: \[(\frac{1}{6},\frac{-1}{3}) and (\frac{5}{6},3)\]

OpenStudy (anonymous):

Honestly, you dont have to help, no one seems to want to help tonight. I am completely lost with these questions, and all everyone can ask is more questions that I do not know the answer to

OpenStudy (anonymous):

Nah, we are going to get this. I want you to learn, but if I just give you the answer, you won't be able to think about how to do it.

OpenStudy (anonymous):

So, point one is: (1/6,-1/3)

OpenStudy (anonymous):

What is the x and y of point 1?

OpenStudy (anonymous):

(x,y)

OpenStudy (anonymous):

(x,y) is the notation of coordinates

OpenStudy (anonymous):

x1 is 1/6 and y1 is -1/3

OpenStudy (anonymous):

Perfect!

OpenStudy (anonymous):

omg, I feel like an idiot

OpenStudy (anonymous):

So what is our x2 and y2?

OpenStudy (anonymous):

Not at all! You got the answer correct so far! We'll just do it one step at a time. You'd only be an idiot if you gave up and said "I'm never going to learn this stuff. Bye." :P

OpenStudy (anonymous):

5/6 and 3

OpenStudy (anonymous):

Exactly. Now do y1 - y2 (just subtract your y2 from y1)

OpenStudy (anonymous):

so y1?

OpenStudy (anonymous):

is anyone there?

OpenStudy (anonymous):

Remember: if the number comes after the letter, it means it's only a 'subscript'. While you could do 2y - 1y = 1y, you have to remember that y1 and y2 are not the same thing. The y1 means "my first y", and y2 means "my second y". You solved what each one means earlier.

OpenStudy (anonymous):

so just 1?

OpenStudy (anonymous):

No, remember that y1 is referring to a point. What point did we say that y1 is?

OpenStudy (anonymous):

I dont remember